# Synthesizing Sentiment-Controlled Feedback For Multimodal Text and Image Data

### v3

@@ Abstract.
The ability to generate sentiment-controlled feedback in response to multimodal inputs comprising text and images addresses a critical gap in human-computer interaction. This capability allows systems to provide empathetic, accurate, and engaging responses, with useful applications in education, healthcare, marketing, and customer service. To this end, we have constructed a large-scale Controllable Multimodal Feedback Synthesis (CMFeed) dataset and propose a controllable feedback synthesis system. The system features an encoder, decoder, and controllability block for textual and visual inputs. It extracts features using a transformer and Faster R-CNN networks, combining them to generate feedback. The CMFeed dataset includes images, texts, reactions to the posts, human comments with relevance scores, and reactions to these comments. These reactions train the model to produce feedback with specified sentiments, achieving a sentiment classification accuracy of 77.23%, which is 18.82% higher than the accuracy without controllability. The system also incorporates a similarity module for assessing feedback relevance through rank-based metrics and an interpretability technique to analyze the contributions of textual and visual features during feedback generation. Access to the CMFeed dataset and the system’s code is available at github.com/MIntelligence-Group/CMFeed.
Affective Computing, Controllability, Interpretability, Multimodal Analysis.
copyright: acmcopyrightjournalyear: 2024doi: XXXXXXX.XXXXXXXjournal: TOMMjournalvolume: 1journalnumber: 1article: 1publicationmonth: 10ccs: Computing methodologies Natural language generationccs: Information systems Sentiment analysisccs: Information systems Multimedia and multimodal retrievalccs: Human-centered computingccs: Computing methodologies Machine learning
1. Introduction
@@
The process of multimodal feedback synthesis involves generating responses to multimodal inputs in a way that mimics human spontaneous reaction (Kumar et al., 2023). Controlling sentiments in feedback, a capability inherent to humans, remains a challenge for machines (Makiuchi et al., 2021). The ability to control sentiments in feedback synthesis facilitates more empathetic responses in healthcare, accurate marketing insights, and engaging educational content while enabling systems to predict patients’ mental states, assess product responses, analyze social behaviors, and gauge user engagement in advertisements (Gallo et al., 2020; Blikstein and Worsley, 2016). The controllability of these systems allows them to be customized for individual users, enhancing personalization (Muszynski et al., 2019).
@@
The need for the Controllable Multimodal Feedback Synthesis (CMFeed) dataset arises from the requirement of a dataset containing human-generated feedback in addition to multimodal inputs. The CMFeed dataset has been created by crawling Facebook news articles and includes input images, texts, human comments, comments’ metadata (such as likes, shares, reactions, and relevance scores), and sentiment labels. Unlike traditional sentiment-controlled text generation systems that do not utilize human comments, systems developed using the CMFeed dataset can be distinctively trained on human-generated comments to learn human-like spontaneity and contextual diversity. This enables the generation of ‘opinions’ rather than just ‘knowledge’ or ‘facts,’ which the proposed task focuses on. Our approach uniquely allows for the generation of controlled opinions.
@@
Based on the CMFeed dataset, a novel task of controllable feedback synthesis for input images and text has been defined, and a feedback synthesis system has been proposed to generate sentiment-controlled feedback. It includes two networks for textual and visual modalities, each with an encoder, a decoder, and a control layer. The encoders use a text transformer (Vaswani et al., 2017) and a Faster R-CNN model (Ren et al., 2015) to extract features, which the decoder then combines for feedback generation. The control layer, positioned after the decoder, selectively activates or deactivates neurons to align with positive or negative sentiments as required. The system also incorporates a similarity module to ensure feedback aligns with the input context. To establish benchmarks for this new task, we compared our system against several baselines using different fusion methods and encoding strategies. The proposed system achieved a sentiment classification accuracy of 77.23%, significantly higher than the baselines. The major contributions of this paper are summarized as follows:
@@
  • A new dataset, CMFeed, has been constructed containing text, images, corresponding comments, number of likes, shares, and sentiment class. This dataset facilitates training multimodal feedback synthesis models to generate sentiment-controlled feedback for input containing images and text.
@@
  • An interpretability technique, K-Average Additive exPlanation (KAAP) has been incorporated that analyzes the contributions of textual and visual features during feedback generation.
@@
  • A novel controllability module has been introduced, enabling sentiment regulation in the generated feedback. It selectively activates or deactivates neurons to ensure alignment of the feedback with the desired sentiment.
The remainder of the paper is organized as follows: the existing related works are surveyed in Section 2. Section 3 describes the dataset construction process and the proposed system. The experiments and results have been discussed in Section 4 and Section 4.4 respectively, and Section 5 outlines the discussions, conclusions and future directions.
2. Related Works
This section explores the related works to the proposed task of controllable feedback synthesis, including multimodal summarization, visual question answering, dialogue generation, and sentiment-aware text generation.
2.1. Multimodal Summarization
Multimodal summarization combines the inputs from various modalities to produce comprehensive summaries (Zhu et al., 2023). In this direction, Chen and Zhuge (Chen and Zhuge, 2018) used recurrent networks to align text with corresponding images. In another work, Zhu et al. (Zhu et al., 2018) focused on extractive summarization, selecting relevant images and texts using a pointer generator network. Their work was extended by Zellers et al. (Zellers et al., 2020) for video summarization, using self-attention mechanisms to choose relevant video frames based on semantic content. Shang et al. (Shang et al., 2021) emphasized timing in video summarization with time-aware transformers. In other studies, Zhao et al. (Zhao et al., 2021) explored audiovisual video summarization including sensory input and Xie et al. (Xie et al., 2022) used visual aesthetics while generating multimodal summaries.
Despite the advancements mentioned above, multimodal summarization faces the challenge of modality bias (Page Fortin and Chaib-draa, 2019; Zhu et al., 2020), and it has not been explored for affect synthesis. Moreover, multimodal summarization approaches do not incorporate human comments in their model training. In contrast, the proposed system is trained on human-generated comments alongside text and image inputs to generate sentiment-controlled feedback. Furthermore, unlike multimodal summarization, which condenses the given information, controllable feedback synthesis produces feedback that aligns with sentiments and fits the context.
2.2. Visual Question Answering
Visual Question Answering (VQA) combines visual perception with interactive question answering (Antol et al., 2015). Chen et al. (Chen et al., 2023) explored controlled generation versus standard answering methods in VQA, while Wang et al. (Wang et al., 2017) focused on fact-based control to improve response accuracy. Cascante et al. (Cascante-Bonilla et al., 2022) introduced an innovative approach using synthetic data, including 3D and physics simulations, to expand VQA’s scope and enhance model adaptability. Furthermore, Jin et al. (Jin et al., 2021) proposed a text-based VQA system that explored the correlation between the text in the images and the visual scene. In another work, Guo et al. (Guo et al., 2021) developed a universal quaternion hypergraph network for advanced multimodal video question answering.
Contributions like Wu et al.’s (Wu et al., 2023) memory-aware control in community question answering and Lehmann et al.’s (Lehmann et al., 2023) use of language models for knowledge graph question answering illustrate VQA’s diverse applications. Furthermore, Ishmam et al. (Ishmam et al., 2024) discuss the contemporary shift towards vision language pre-training techniques and highlight the challenges in integrating traditional VQA approaches with modern VLP methods. The developments in VQA, however, stand in contrast to the proposed task of sentiment-controlled feedback synthesis. While VQA focuses on generating accurate responses to questions, sentiment-controlled feedback synthesis is centered on creating responses to multimodal stimuli (images and text) that are contextually relevant.
2.3. Dialogue Generation
In dialogue generation, particularly in visual dialogue (VisDial), computational models are designed to engage in dialogues about the visual content present in images. In this direction, Kang et al. (Kang et al., 2019) developed a dual Attention Network that utilized the understanding of visual references in dialogues. This approach was further advanced by Jiang et al. (Jiang et al., 2020), who implemented a region-based graph attention network for enhanced image-related questioning. In other works, Xu et al. (Xu et al., 2018) and Zhao et al. (Zhao et al., 2017) worked on modeling the conversations utilizing conditional variational autoencoders for diverse dialogue scenarios.
@@ 2.3. Dialogue Generation
Recent developments in VisDial include the work of Chen et al. (Chen et al., 2022), who applied contrastive learning for better cross-modal comprehension in visual dialogue. Contributions from Wang et al. (Wang et al., 2022) and Kang et al. (Kang et al., 2023) involved enhanced dialogue generation through generative self-training. Liu et al. (Liu et al., 2022) and Liu et al. (Liu et al., 2023) also contributed to closed-loop reasoning and counterfactual visual dialogue for unbiased knowledge training. Despite the aforementioned developments, existing VisDial methods do not generate sentiment-controlled feedback. The proposed system stands apart as it synthesizes feedback that is not only contextually relevant but is also tailored to the sentiment of the conversation.
2.4. Sentiment-Aware Conversation Agents
In the context of sentiment-aware conversational agents, Das et al. (Das et al., 2017a) laid the groundwork by using multiple generators and a multi-class discriminator to create text with diverse sentiment tones. Building upon this, Shi et al. (Shi and Yu, 2018) introduced a framework leveraging user sentiments to guide dialogue actions. In another work, Kong et al. (Kong et al., 2019) explored sentiment-constrained dialogue generation using generative adversarial networks. In recent developments, Firdaus et al. (Firdaus et al., 2020) presented a system that integrates sentiments into conversational agents. Their SEPRG model complemented it (Firdaus et al., 2021), which focuses on sentiment-aware, sentiment-controlled personalized response generation. In other works, Hu et al. (Hu et al., 2022) developed a speech sentiment-aware agent for empathetic responses, and Saha et al. (Saha et al., 2022) integrated sentiment-aware mechanisms into conversation generation. Unlike the aforementioned works that primarily focus on recognizing the users’ sentiments and using the same to influence dialogue actions, the proposed system generates (as opposed to recognizing) the feedbacks with the desired sentiment tone.
2.5. Sentiment-Controlled Text Generation
Recent natural text generation research trends include sentiment-controlled text generation. In this context, Zhou et al. (Zhou et al., 2018) introduced the emotional chatting machine, employing a seq2seq framework with internal and external memories for managing emotional states. In another work, Wang and Wan (Wang and Wan, 2018) generated sentimental texts using mixture adversarial networks. Huang et al. (Huang et al., 2018) contributed to automatic dialogue generation with a focus on expressed emotions. Building on these, Zhong et al. (Zhong et al., 2019) introduced an affective attention mechanism with a weighted cross-entropy loss for generating affective dialogue. Extending the scope of sentiment control into other domains, Wu et al. (Wu et al., 2024) developed Music ControlNet, a diffusion-based model that manipulates time-varying controls such as melody, rhythm, and dynamics in music generation, paralleling the challenges in sentiment-controlled text generation. Similarly, Yang et al. (Yang et al., 2024) enhanced text-to-image models by employing glyph information to precisely control text placement and size in images, demonstrating advanced customization capabilities in multimodal interactions.
The sentiment-controlled text generation methods do not use human comments to train their models. On the contrary, the proposed system captures the wider sentiment context from multimodal (text and images) inputs. Moreover, it is trained on human-generated comments aside from text and images to learn the contextual diversity from the comments. Furthermore, our work generates ‘opinions’ in contrast to generation of ‘knowledge’ or ‘facts’ by the above mentioned tasks. Unlike facts, the opinions can be controlled in a similar way how humans do. This area has not been addressed by the tasks mentioned earlier, while our work actively contributes towards it.
@@ 2.5. Sentiment-Controlled Text Generation
Table 1. Related datasets. ‘HC’: Human Comments, ‘CM’: Comments’ Metadata, ‘V’: Visual, ‘A’: Audio, ‘T’: Textual, ‘MMSum’: Multimodal Summarization, ‘VQA’: Vis Qu Ans, ‘VisDial’: Vis Dialogue, ‘SATextGen’: Sentiment-Aware Text Generation, ‘N/A’: Not Applicable, ‘-’: Unavailable.
@@
MMSum VMSMO (Zellers et al., 2020) 2020 184920 documents 70 participants V, T
MSMO (Zhu et al., 2018) 2018 314581 documents 10 students V, T
MVSA-Single (Niu et al., 2016) 2016 4869 tweets N/A V, T
MVSA-Multiple (Niu et al., 2016) 2016 4869 tweets N/A V, T
VQA DocVQA (Mathew et al., 2020) 2020 50K QA pairs - T
OK VQA (Marino et al., 2019) 2019 150K QA pairs 5 MTurk workers V, T
VQA (Goyal et al., 2017) 2017 1.1M QA pairs 215 MTurk workers V, T
VideoQA (Xu et al., 2017) 2017 243K QA pairs - V, T
VisDial InfoVisDial (Wen et al., 2023) 2023 79535 dialogues 35 annotators V, T
CLEVR-Dialog (Kottur et al., 2019) 2019 4.25M dialogues - V, T
VisDial (Das et al., 2017b) 2017 1.23M QA pairs 200 annotators V, T
Visual Madlibs (Yu et al., 2015) 2015 397675 dialogues 1 quality checker V, T
SATextGen SEPRG (Firdaus et al., 2021) 2021 64356 conversations 500 samplers T
EMOTyDA (Saha et al., 2020) 2020 19365 videos 10 participants V, T
ESTC (Zhou et al., 2018) 2018 4308211 conversations - T
STC (Shang et al., 2015) 2015 4.4M conversations - T
The datasets used for the aforementioned methods are summarized in Table 1. While a few include human-generated summaries or answers, none offer the metadata (such as relevance, reactions, and sentiment labels) necessary for training feedback synthesis systems. The absence of human comments and metadata in existing datasets hinders the development of sentiment-controlled feedback synthesis systems (Kumar et al., 2023; Wu et al., 2019). In contrast, CMFeed addresses this gap by supporting the development of models that can generate opinions mimicking human conversational dynamics, a feature not covered by existing tasks. This capability enables CMFeed to create more personalized and contextually aware multimodal interactions, allowing sentiment to reflect the natural ebb and flow of human emotions. In contrast to the latest large language models (LLMs) such as Generative Pre-Trained Transformer (GPT) (Radford et al., 2018), which primarily excel in summarization and question answering through knowledge retrieval, the CMFeed dataset is designed to interpret multimodal inputs and produce novel, sentiment-controlled feedback that mimics human interaction, thereby marking a distinct contribution from conventional LLMs that rely on existing knowledge. Further, LLMs are prone to mistakes and hallucinations (Farquhar et al., 2024; Huang et al., 2023). They are auto-regressive and black-box. Autoregressive means if the first few words are wrong, it may lead to incorrect generation of subsequent words. The internal mechanisms, such as which modality influences the response, are generally not transparent. By incorporating the proposed interpretability mechanism detailed in Sections 3.3.7 and 4.4.5, we have demonstrated how metadata and multimodal features influence responses and how sentiments are learned. In fact, the proposed concepts of controllability and interpretability can inspire new ideas for the LLMs and other related tasks such as VisDial.
3. Proposed Methodology
3.1. Dataset Construction
3.1.1. Data collection
@@
The CMFeed dataset has been compiled by crawling news articles from Sky News, NYDaily, FoxNews, and BBC News through Facebook posts. The data collection process utilized the NLTK (Bird et al., [n. d.]) & newspaper3k (Ou-Yang, [n. d.]) libraries and it was conducted in compliance with Facebook’s terms and conditions (Meta, [n. d.]), ensuring adherence to all legal and ethical standards.
@@
Choice of Facebook for Dataset Construction: Facebook was chosen for data collection due to its unique provision of metadata such as news article links, post shares, post reactions, comment likes, comment rank, comment reaction rank, and relevance scores, which are not available on other platforms. With $3.07$ billion monthly users, it is the most used social media platform, compared to Twitter’s 550 million and Reddit’s 500 million (Wikipedia, 2024). Broad popularity is maintained across various age groups, with at least 58% usage among ages 18-29, 30-49, 50-64, and above 65, significantly surpassing the 6% for Twitter and 3% for Reddit (Social, 2024). This trend is consistent across demographics such as gender, race, ethnicity, income, education, community, and political affiliation (Center, 2024). The male-to-female user ratio is reported as 56.3% to 43.7% on Facebook, compared to Twitter’s 66.72% to 23.28%; such data are not reported by Reddit (Facebook, 2024a). These characteristics make Facebook an ideal platform for the collection of a diverse and representative dataset.
@@
Choice of News Handles: Four news handles—BBC News, Sky News, Fox News, and NY Daily News—were selected to ensure diversity and comprehensive regional coverage. These news outlets were chosen for their distinct editorial perspectives and regional focuses: global coverage and centrist views are known from BBC News, targeted content and center/right leanings in the UK, EU, and US are offered by Sky News, right-leaning content in the US is recognized from Fox News, and left-leaning coverage in New York is provided by NY Daily News. A broad spectrum of political discourse and audience engagement is ensured by this selection.
3.1.2. Preprocessing

### v4

@@ Abstract
The ability to generate sentiment-controlled feedback in response to multimodal inputs comprising text and images addresses a critical gap in human-computer interaction. This capability allows systems to provide empathetic, accurate, and engaging responses, with useful applications in education, healthcare, marketing, and customer service. To this end, we have constructed a large-scale Controllable Multimodal Feedback Synthesis (CMFeed) dataset and proposed a controllable feedback synthesis system. The system features an encoder, decoder, and controllability block for textual and visual inputs. It extracts features using a transformer and a Faster R-CNN network, combining them to generate feedback. The CMFeed dataset includes images, texts, reactions to the posts, human comments with relevance scores, and reactions to these comments. These reactions train the model to produce feedback with specified sentiments, achieving a sentiment classification accuracy of 77.23%, which is 18.82% higher than the accuracy without controllability. Access to the CMFeed dataset and the system’s code is available at github.com/MIntelligence-Group/CMFeed.
I Introduction
@@
The process of multimodal feedback synthesis involves generating responses to multimodal inputs in a way that mimics human spontaneous reactions [1]. Controlling sentiments in feedback, a capability inherent to humans, remains a challenge for machines [2]. The ability to control sentiments in feedback synthesis facilitates more empathetic responses in healthcare, accurate marketing insights, and engaging educational content while enabling systems to predict patients’ mental states, assess product responses, analyze social behaviors, and gauge user engagement in advertisements [3, 4]. The controllability of these systems allows them to be customized for individual users, enhancing personalization [5]. Our goal is to advance human-centered, controllable interaction by explicitly regulating sentiment and exposing interpretable control signals, strengthening user understanding and trust in the loop.
@@
The need for the Controllable Multimodal Feedback Synthesis (CMFeed) dataset arises from the requirement of a dataset containing human-generated feedback in addition to multimodal inputs. The CMFeed dataset has been created by crawling Facebook news articles and includes input images, texts, human comments, comments’ metadata (such as likes, shares, reactions, and relevance scores), and sentiment labels. Unlike traditional sentiment-controlled text generation systems that do not utilize human comments, systems developed using the CMFeed dataset can be distinctively trained on human-generated comments to learn human-like spontaneity and contextual diversity. This enables the generation of ‘opinions’ rather than just ‘knowledge’ or ‘facts,’ which the proposed task focuses on.
@@
In this work, a novel task of controllable feedback synthesis for input images and text has been defined, and a feedback synthesis system has been proposed to generate sentiment-controlled feedback. It includes two networks for textual and visual modalities, each with an encoder, a decoder, and a control layer. The encoders use a text transformer [6] and a Faster R-CNN model [7] to extract features, which the decoder then combines for feedback generation. The control layer, positioned after the decoder, selectively activates or deactivates neurons to align with positive or negative sentiments as required. The system also incorporates a similarity module to ensure feedback aligns with the input context. The proposed system achieved a sentiment classification accuracy of 77.23% on the CMFeed dataset, significantly higher than the baselines. The major contributions of this paper are summarized as follows:
@@
  • A new dataset, CMFeed, has been constructed containing text, images, corresponding comments, number of likes, shares, and sentiment class. It enables training models to generate sentiment-controlled feedback from image–text inputs.
@@
  • A novel controllability module has been proposed that selectively activates or deactivates neurons to align the feedback with the desired sentiment.
  • An interpretability technique, K-Average Additive exPlanation (KAAP) has been incorporated to analyze of textual and visual features’ contribution during feedback generation.
II Related Works
II-A Multimodal Summarization
Multimodal summarization combines inputs from various modalities to produce comprehensive summaries [8]. Zhu et al. [9] focused on extractive summarization with a pointer generator; Li et al. [10] extended this to video with self-attention for frame selection. Other studies include Zhao et al. [11] on audio-visual summarization and Xie et al. [12] on visual aesthetics. Despite advances, multimodal summarization faces modality bias [13, 14], is not explored for affect synthesis, and does not incorporate human comments in training. In contrast, our system is trained on human-generated comments alongside text and images to generate sentiment-controlled feedback; rather than condensing information, it produces feedback aligned with sentiments and context.
II-B Visual Question Answering
Visual Question Answering (VQA) combines visual perception with interactive question answering [15]. Chen et al. [16] explored controlled generation versus standard answering; Wang et al. [17] focused on fact-based control. Cascante et al. [18] introduced synthetic data to expand scope, and Guo et al. [19] proposed a quaternion hypergraph network for multimodal video QA. Wu et al. [20] added memory-aware control in community QA, and Lehmann et al. [21] used language models for knowledge graph QA, illustrating diverse applications and the shift toward vision and language pre-training. Unlike our task of sentiment-controlled feedback synthesis, which creates contextually relevant responses to multimodal stimuli (images and text), VQA focuses on accurate answers to questions.
II-C Dialogue Generation
@@
In dialogue generation, particularly visual dialogue (VisDial), computational models are designed to converse about visual content [22]. Kang et al. [23] developed a dual Attention Network that utilized the understanding of visual references in dialogues. Jiang et al. [24] implemented a region-based graph attention network for enhanced image-related questioning. Contributions from Wang et al. [25] and Kang et al. [26] enhanced dialogue generation through generative self-training. Liu et al. [27] contributed to closed-loop reasoning and counterfactual visual dialogue for unbiased knowledge training. Despite these advancements, existing VisDial methods do not generate sentiment-controlled feedback. The proposed system stands apart as it synthesizes feedback that is contextually relevant and tailored to the sentiment of the conversation.
II-D Sentiment-Aware and Sentiment-Controlled Generation
In sentiment-aware conversational agents, Shi et al. [28] leverage user sentiments to guide dialogue actions; variational models predict emotional reactions to utterances for contextual guidance [29]; Firdaus et al. [30] integrate sentiments for sentiment-controlled personalized responses; Hu et al. [31] design a speech sentiment-aware agent. Controllability has also extended to other modalities, e.g., Music ControlNet for melody/rhythm [32] and glyph-conditioned control for text-to-image precision [33]. However, sentiment-controlled text generation does not use human comments. By contrast, our system captures wider sentiment context from multimodal (text and images) inputs and is trained on human-generated comments to learn contextual diversity. Our work generates ‘opinions’ rather than ‘knowledge’ or ‘facts’; unlike facts, opinions can be controlled similarly to humans. This area has not been addressed by earlier tasks that our work advances.
@@ II-D Sentiment-Aware and Sentiment-Controlled Generation
TABLE I: Summary of the existing related datasets. Here, ‘HC’: Human Comments, ‘CM’: Comments’ Metadata, ‘V’: Visual, ‘A’: Audio, ‘T’: Textual, ‘MMSum’: Multimodal Summarization, ‘VQA’: Visual Question Answering, ‘VisDial’: Visual Dialogue Generation, ‘SATextGen’: Sentiment-Aware Text Generation, and ‘N/A’: Not Applicable, ‘-’: Unavailable.
@@
MMSum
VMSMO [10] 2020 184920 documents 70 participants V, T
MSMO [9] 2018 314581 documents 10 students V, T
MVSA-Single [34] 2016 4869 tweets N/A V, T
MVSA-Multiple [34] 2016 4869 tweets N/A V, T
VQA
DocVQA [35] 2020 50K QA pairs - T
OK VQA [36] 2019 150K QA pairs 5 MTurk workers V, T
VQA [37] 2017 1.1M QA pairs 215 MTurk workers V, T
VideoQA [38] 2017 243K QA pairs - V, T
VisDial
InfoVisDial [39] 2023 79535 dialogues 35 annotators V, T
CLEVR-Dialog [40] 2019 4.25M dialogues - V, T
VisDial [41] 2017 1.23M QA pairs 200 annotators V, T
Visual Madlibs [42] 2015 397675 dialogues 1 quality checker V, T
SATextGen
SEPRG [43] 2021 64356 conversations 500 samplers T
EMOTyDA [44] 2020 19365 videos 10 participants V, T
ESTC [45] 2018 4308211 conversations - T
STC [46] 2015 4.4M conversations - T
Existing datasets are summarized in Table I. While some include human-generated summaries, none provide metadata such as relevance, reactions, or sentiment labels for feedback synthesis systems. The lack of human comments and metadata limits sentiment-controlled feedback synthesis [1, 47]. CMFeed fills this gap by enabling models to generate opinions that mimic human conversational dynamics, supporting context-aware multimodal interactions. Unlike large language models (LLMs) such as GPT [48], which focus on summarization and question answering, CMFeed interprets multimodal inputs to produce sentiment-controlled feedback. LLMs are prone to errors and hallucinations [49, 50], are autoregressive, and obscure modality influence. Our interpretability mechanism shows how metadata and multimodal features shape responses and sentiment, and these notions of controllability and interpretability can inform future LLM and VisDial research.
III Proposed Methodology
III-A Dataset Construction
III-A1 Data Collection
@@
The CMFeed dataset has been compiled by crawling news articles from Sky News, NYDaily, FoxNews, and BBC News through Facebook posts. The data collection process utilized the NLTK [51] & newspaper3k [52] libraries and it was conducted in compliance with Facebook’s terms and conditions [53], ensuring legal and ethical standards.
@@
Choice of Facebook for Dataset Construction: Facebook was chosen for data collection due to its unique provision of metadata, i.e., news article links, post shares, post reactions, comment likes, comment rank, comment reaction rank, and relevance scores are unavailable on other platforms. With $3.07$ billion monthly users, it is the most used social media platform, compared to Twitter’s 550 million and Reddit’s 500 million [54]. Popularity holds across age groups, with at least 58% usage among ages 18–29, 30–49, 50–64, and 65+, versus 6% for Twitter and 3% for Reddit [55]. This trend is consistent across gender, race/ethnicity, income, education, community, and political affiliation [56]. The male-to-female user ratio is 56.3% to 43.7% on Facebook, compared to Twitter’s 66.72% to 23.28%; such data are not reported by Reddit [57].
@@
Choice of News Handles: Four news handles (BBC News, Sky News, Fox News, and NY Daily News) were selected to ensure diversity and comprehensive regional coverage. These news outlets were chosen for their distinct editorial perspectives and regional focuses: global coverage and centrist views are known from BBC News, targeted content and center/right leanings in the UK, EU, and US are offered by Sky News, right-leaning content in the US is recognized from Fox News, and left-leaning coverage in New York is provided by NY Daily News. A broad spectrum of political discourse and audience engagement is ensured by this selection.
III-A2 Preprocessing